# Closing the Gap on the Sample Complexity of 1-Identification

### v1

@@ 1.3 Main Contribution
In this paper, we develop novel upper and lower bounds for the 1-identification problem. Regarding the lower bound, we conclude
@@
For any $\delta$-PAC alg and any reward vector $\{\mu_{a}\}_{a=1}^{K}\in[0,1]^{K}$ satisfying that $\mu_{1}\geq\cdots\geq\mu_{m}>\mu_{0}\geq\mu_{m+1}\geq\cdots\geq\mu_{K}$, any tolerance level $\delta$ such that $\delta<\min\{\frac{1}{10^{-8}},\frac{1}{64m^{2}}\}$, we can find a permutation $\sigma:[K]\rightarrow[K]$ and define a positive instance $\nu$ with $r_{\sigma(a)}^{\nu}=\mu_{a},a\in[K]$ with unit-variance Gaussian reward, such that $\mathbb{E}_{\text{alg},\nu}\tau\geq\Omega\left(\min\limits_{1\leq j\leq m}\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}+\frac{H(j)}{\log^{2}(m+1)}\right)$.
Regarding the upper bound, we merge and adapt the idea in Katz-Samuels and Jamieson (2020) and Li and Cheung (2025) and design a new algorithm Parallel Sequential Exploration Exploitation on Brackets(Algorithm 1, PSEEB). We can prove the following two theorems.
@@
$\mathbb{E}_{\nu}\tau\leq\begin{cases}O\left(\log K\min\limits_{1\leq j\leq m}\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}+\log^{3}K(\log(4K)+\log_{2}\frac{1}{\Delta_{j,0}^{2}})H(j)\right)&\nu\text{ is positive}\\ O\left((\log K)H_{1}^{\text{neg}}\log\frac{H_{1}^{\text{neg}}}{\delta}\right)&\nu\text{ is negative}\end{cases}$
combining Theorem 2 and 3, we know algorithm 1 is $(\Delta,\delta)$-PAC. What’s more, the gap between Theorem 1 and 3 is up to a polynomial of $(\log K,\log\{\frac{1}{\Delta_{0,a}^{2}}\}_{a=1}^{m})$, suggesting that our proposed algorithm 1 is nearly optimal for all 1-identification instances, and $\min_{1\leq j\leq m}\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}+H(j)$ is an appropriate complexity measure. Compared to history literature, our proposed Algorithm 1 complement one of the open direction left in Li and Cheung (2025), by deriving a new upper bound for $\mathbb{E}\tau$ when there are multiple qualified arms.
@@ 1.4 Literature Review on Other Related Works
If we consider the given threshold $\mu_{0}$ as an arm with deterministic reward, we can address 1-identification by solving a Best Arm Identification(BAI) problem Even-Dar et al. (2002), Gabillon et al. (2012), Jamieson et al. (2013), Kalyanakrishnan et al. (2012), Karnin et al. (2013), Jamieson and Nowak (2014), Chen and Li (2015). However, this is an inefficient solution, suggested by the asymptotic result in Degenne and Koolen (2019), Garivier and Kaufmann (2016) and non-asymptotic result in Li and Cheung (2025), Kano et al. (2017). The reason is 1-identification doesn’t care about an exact comparison between arms with high confidence level, and we should not pull arms as frequently as BAI to distinguish two arms.
Besides 1-identification, there are other variants of pure exploration with multiple answers, as mentioned by Degenne and Koolen (2019). The agent needs to find all the arms Mason et al. (2020) in $\{a:\mu_{a}\geq\mu_{1}-\epsilon\}$ or only one arm Even-Dar et al. (2002), Gabillon et al. (2012), Kalyanakrishnan et al. (2012), Kaufmann and Kalyanakrishnan (2013), Katz-Samuels and Jamieson (2020). As discussed in Katz-Samuels and Jamieson (2020), if the gap $\mu_{1},\mu_{0}$ are known and $\mu_{1}>\mu_{0}$, by taking $\epsilon=\Delta_{0,1}$, the answer set of $\epsilon$-Good Arm Identification and 1-identification are the same. We can adopt same pulling strategy and recommendation rule. But it is unclear about the relationship if there are unknown. What’s more, Locatelli et al. (2016), Mukherjee et al. (2017) work on Thresholding Bandits, which aims to output a set of arms whose reward are above a given $\mu_{0}$, within fixed budget. The optimal set might not be unique under their proposed problem setting.
In this paper, we adopt the idea of parallel algorithm copies, which also exists in the history literature. Chen and Li (2015), Chen et al. (2017), Katz-Samuels and Jamieson (2020) all require a main algorithm to manage different algorithm copies. Katz-Samuels and Jamieson (2020) allocate $2^{\ell}$ time rounds to a copy each time with increasing index $\ell$. Chen and Li (2015), Chen et al. (2017) maintain an overall round index $t\in\mathbb{N}$ and further assign algorithm index to each copy. They execute an algorithm copy only when its algorithm index divide $t$. In our work, we apply round-robin procedure to execute each algorithm copy.
2 Lower Bound of Sample Complexity, for a Positive Instance
@@
In this section, we continue to work on $\{\mu_{a}\}_{a=1}^{K}$ such that $\mu_{1}\geq\mu_{2}\geq\cdots\geq\mu_{m}>\mu_{0}\geq\mu_{m+1}\geq\cdots\geq\mu_{K}$ and instance $\nu$ with $r_{a}^{\nu}=\mu_{a}$. We assume each reward distribution follows a unit-variance Gaussian distribution in our lower bound analysis. For an instance $\nu^{\prime}$ and a permutation $\sigma:[K]\rightarrow[K]$, represented as a bijection from $[K]$ to $[K]$, we define the instance $\sigma(\nu^{\prime})$ as $r^{\sigma(\nu^{\prime})}_{\sigma(a)}=r_{a}^{\nu^{\prime}}$.
We first introduce the definition of a symmetric algorithm.
Definition 3.
@@
An algorithm $\widetilde{\text{alg}}$ is said to be symmetric if the following statement holds. For any instance $\nu$ with reward vector $\{\mu_{a}\}_{a=1}^{K}$, any sequence $\{a_{t}\}_{t=1}^{T}\in[K]^{T}$, natural number $T\geq 1$, and two different permutations $\sigma^{\prime},\sigma^{\prime\prime}$, the algorithm $\widetilde{\text{alg}}$ satisfies
$\Pr_{\sigma^{\prime}(\nu),\tilde{\text{alg}}}(\forall t\in[T],A_{t}=\sigma^{\prime}(a_{t}),\tau>T)=\Pr_{\sigma^{\prime\prime}(\nu),\tilde{\text{alg}}}(\forall t\in[T],A_{t}=\sigma^{\prime\prime}(a_{t}),\tau>T).$
@@
Intuitively, under a symmetric algorithm $\widetilde{\text{alg}}$, the reward distribution of the pulled arm $A_{t}$ at time $t$ is invariant under any permutation on the arms’ indexes. In particular, for a symmetric algorithm $\widetilde{\text{alg}}$, we have $\mathbb{E}_{\sigma^{\prime}(\nu),\tilde{\text{alg}}}\tau=\mathbb{E}_{\sigma^{\prime\prime}(\nu),\tilde{\text{alg}}}\tau$ for any two permutations $\sigma^{\prime},\sigma^{\prime\prime}$. A similar definition also appears in Definition 12 of Simchowitz et al. (2017), which further requires the distribution of output $\hat{a}$ is also invariant regarding the distribution. This extra requirement is unnecessary in our analysis. By the following lemma, it suffices to consider only symmetric algorithms for establishing our lower bound.
@@
Denote $S$ as the set of all permutations of $[K]$. Given any $\delta$-PAC algorithm alg, we can always find a $\delta$-PAC symmetric algorithm $\widetilde{\text{alg}}$, such that for any instance $\nu^{\prime}$, we have $\max\limits_{\sigma\in S}\mathbb{E}_{\text{alg},\sigma(\nu^{\prime})}\tau\geq\mathbb{E}_{\tilde{\text{alg}},\nu^{\prime}}\tau$.
Full proof of Lemma 4 is in Appendix B.1. We can shuffle all arms before applying any alg to create a symmetric algorithm $\widetilde{\text{alg}}$. Given Lemma 4, to prove Theorem 1, we suffice to prove the following Theorem.
@@
Suppose $\delta<\min\{10^{-8},1/(64m^{2})\}$. For any $\delta$-PAC symmetric algorithm, and for any positive instance $\nu$ with $m\geq 1$ eligible arms, we have
$\mathbb{E}\tau\geq\Omega\left(\min_{1\leq j\leq m}\frac{\log(1/\delta)}{\Delta_{j,0}^{2}}+\frac{1/j}{\log^{2}(m+1)}\sum_{a=1}^{K}\frac{1}{\max\{\Delta_{a,0}^{2},\Delta_{a,j}^{2}\}}\right).$
The assumption $\delta<\min\{10^{-8},1/(64m^{2})\}$ is to ensure $\delta\cdot m^{2}$ is upper bounded by a non-trivial absolute constant. The upper bound eliminates the possibility of being $\delta$-PAC by outputting a uniformly chosen arm without any arm pulls.
@@
In what follows, we present the main steps to prove Theorem 5. The full proof is in Appendix B.2, which consists of the statements and proofs of auxiliary lemmas. The main idea is to establish Lemma 6,
@@
Assume $\delta<10^{-8}$, $\frac{1}{8m}\geq\sqrt{\delta}$, we can assert $\mathbb{E}_{\nu}\tau$ is lower bounded by the $\frac{1}{3200}$ times optimal value of the following optimization problem
@@
If Lemma 6 holds, we first prove the above optimization problem (2) is convex. Then, we turn to construct a feasible solution to the dual problem, and prove the objective value of this feasible solution is lower bounded by $\Omega\left(\min\limits_{1\leq j\leq m}\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}+\frac{H(j)}{\log^{2}(m+1)}\right)$. By the strong duality, we complete the proof of Theorem 5.
Denote event $\mathcal{E}_{j}=\left\{N_{j}(\tau)\geq\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}},\hat{a}=j\right\}$. To prove Lemma 6, we are going to show taking $v=3200\mathbb{E}_{\nu}\tau$, $p_{j}=\Pr_{\nu}(\mathcal{E}_{j})$ for $j\in[m]$ is a feasible solution to (2). The overall idea is to apply the Transportation Equality(Lemma 18 in Kaufmann et al. (2016)) to prove the relationship between $\Pr_{\nu}(\mathcal{E}_{j})$ and $\Pr_{\nu^{\prime}}(\mathcal{E}_{j})$ for some alternative instances $\nu^{\prime}$, further the relationship between $3200\mathbb{E}_{\nu}\tau$ and $\{\Pr_{\nu}(\mathcal{E}_{j})\}_{j=1}^{m}$. We prove the feasibility of $v=3200\mathbb{E}_{\nu}\tau$, $p_{j}=\Pr_{\nu}(\mathcal{E}_{j})$ by going through the constraint (C1), (C2), (C3), (C4) (C5).
For (C1), (C2), they are trivial. As the expected pulling times must be non-negative and the probability must be in the interval $[0,1]$.
@@
$3200\mathbb{E}\tau\geq$ $\sum_{j=1}^{m}\mathbb{E}\tau\mathds{1}(\mathcal{E}_{j})\geq\sum_{j=1}^{m}\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}}\Pr_{\nu}(\mathcal{E}_{j}).$
For (C4), we start by proving the inequality (3).
$\Pr_{\nu}(\tau<+\infty,\hat{a}\in[m])\geq 1-5\delta.$ (3)
To prove (3), we consider two cases, $\{a:\mu_{a}=\mu_{0}\}$ is empty or not. The $\delta$-PAC requirement suggests that (3) must hold for both cases.
Then, we are ready to prove (C4) by contradiction. Denote $\bar{\mathcal{E}_{j}}=\{N_{j_{0}}(\tau)<\frac{\log\frac{1}{\delta}}{100\Delta_{j_{0},0}^{2}},\hat{a}=j_{0}\}$. If $\sum_{j=1}^{m}\Pr_{\nu}(\mathcal{E}_{j})<\frac{1}{2}$, together with (3), we can find a $j_{0}\in[m]$ such that
@@
We consider an alternative instance $\nu^{\prime}_{j_{0}}(\epsilon)$ with $r_{a^{\prime}}^{\nu^{\prime}_{j_{0}}}=\begin{cases}\mu_{0}-\epsilon&a^{\prime}=j_{0}\\ r_{a^{\prime}}^{\nu}&else\end{cases}$. The $\delta$-PAC requirement of algorithm conclude
@@
Combining (5),(4), we have $\log\frac{\Pr_{\nu}\left(\bar{\mathcal{E}}_{j_{0}}\right)}{\Pr_{\nu^{\prime}_{j_{0}}(\epsilon)}(\bar{\mathcal{E}}_{j_{0}})}\geq\frac{1}{2}\log\frac{1}{\delta}$ holds for all $\epsilon>0$. Meanwhile, Lemma 8, which utilizes the Transportation Equality(Lemma 18 in Kaufmann et al. (2016)), suggests that
@@
for each $j\in[m],a\in[K]$, where $T_{j}(a)=\lceil\frac{1}{200\max\{\Delta_{j,0}^{2},\Delta_{a,j}^{2}\}\Big(1+\log\frac{1}{\Pr_{\nu}(\mathcal{E}_{j})}\Big)}\rceil$. If Lemma 11 holds, we can sum up (7) for each $a\in[K]$ and use the fact that $\frac{1}{\max\{\Delta_{j,0}^{2},\Delta_{a,j}^{2}\}}\geq\frac{1}{4\max\{\Delta_{a,0}^{2},\Delta_{a,j}^{2}\}}$, concluding (C5) holds. Then, the remaining work is to prove (7). In the following, we present the sketch idea, which highly relies on the definition of $T_{j}(a)$.
For $n\in\mathbb{N}$, denote
$S_{j,a}^{n}=$ $\{t:N_{j}(t)+N_{a}(t)=n,A_{t+1}\in\{j,a\}\}$
$\tau_{n}^{j}(a)=$ $\min S_{j,a}^{n}$
For any sample path $\{U_{s},A_{s},X_{s}\}_{s=1}^{\tau}$, $S_{j,a}^{n}$ contains at most one element, which is the round that an agent is going to pull arm $a$ or $j$. We suffice to focus on these rounds. In the case that $S_{j,a}^{n}=\emptyset$, we take $\tau_{n}^{j}(a)=+\infty$, which means $\tau_{n}^{j}(a)\geq\tau$. Since $A_{t+1}$ only depends history $H_{t}$, we know $\tau_{n}^{j}(a)$ is a stopping time. Regarding $\tau_{n}^{j}(a)$, we are going to prove two inequalities
$\Pr_{\nu}(\tau_{n}^{j}(a)<\tau)\geq\Pr_{\nu}(\mathcal{E}_{j}),\forall n\leq T_{j}(a).$ (8)
$\Pr_{\nu}(A_{\tau_{n}^{j}(a)+1}=a,\tau_{n}^{j}(a)<\tau)\geq\frac{1}{4}\Pr_{\nu}(\tau_{n}^{j}(a)<\tau),\forall n\leq T_{j}(a).$ (9)
If (8) (9) both hold, from the fact that we only pull arm a at $\tau_{n}^{j}(a)+1$, we can conclude
$\mathbb{E}_{\nu}N_{a}(\tau)\geq$ $\sum_{n=0}^{T_{j}(a)}\Pr_{\nu}(A_{\tau_{n}^{j}(a)+1}=a,\tau_{n}^{j}(a)<+\infty)$
$\geq$ $\sum_{n=0}^{T_{j}(a)}\frac{1}{4}\Pr_{\nu}(\tau_{n}^{j}(a)<\tau)$
$\geq$ $\frac{1}{4}\sum_{n=0}^{T_{j}(a)}\Pr_{\nu}(\mathcal{E}_{j})$
$=$ $\frac{1}{4}T_{j}(a)\Pr_{\nu}(\mathcal{E}_{j}).$
The second step is by (9) and the third step is by (8). Thus, to prove (7), we suffice to show (8) (9) are both correct.

### v2

@@ 1.3 Main Contribution
In this paper, we develop novel upper and lower sample complexity bounds for the 1-identification problem. Our results characterize the sample complexity in both positive and negative instances and close the gap for instances with multiple qualified arms. We first establish a new lower bound for positive instances.
@@
For any $\delta$-PAC alg and any reward vector $\{\mu_{a}\}_{a=1}^{K}\in[0,1]^{K}$ satisfying that $\mu_{1}\geq\cdots\geq\mu_{m}>\mu_{0}\geq\mu_{m+1}\geq\cdots\geq\mu_{K}$, any tolerance level $\delta$ such that $\delta<\min\{10^{-8},\frac{1}{64m^{2}}\}$, we can find a permutation $\sigma:[K]\rightarrow[K]$ and define a positive instance $\nu$ with $r_{\sigma(a)}^{\nu}=\mu_{a},a\in[K]$ with unit-variance Gaussian reward, such that $\mathbb{E}_{\text{alg},\nu}\tau\geq\Omega\left(\min\limits_{1\leq j\leq m}\left\{\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}+\frac{H(j)}{\log^{2}(m+1)}\right\}\right)$.
Theorem 1 is our main contribution, achieving nearly joint optimality on both $\delta$-dependent and $\delta$-independent term. To the best of our knowledge, the “Simulator” framework Simchowitz et al. (2017); Katz-Samuels and Jamieson (2020) was the only known method for deriving $\delta$-independent lower bounds. However, results in Katz-Samuels and Jamieson (2020) are only tight under restricted settings ($\mu_{1}=\dots=\mu_{m}>\mu_{0}$), without joint optimality on both parts. The bottleneck is constructing a “Simulator” that depends on the sampling path and the reward vector. Our proposed method reduces this complex construction to finding a feasible solution for a convex optimization problem, which not only significantly simplifies the proof burden but also yields a tighter, jointly optimal result. We provide sketch proof in Section 2 and full proof in Appendix B.
On the algorithmic side, we combine and refine ideas from Katz-Samuels and Jamieson (2020); Li and Cheung (2025) to propose a new algorithm, Parallel Sequential Exploration–Exploitation on Brackets (Algorithm 1, abbreviated as PSEEB). We establish the following guarantees.
@@
$\mathbb{E}_{\nu}\tau\leq\begin{cases}O\left(\log K\min\limits_{1\leq j\leq m}\left\{\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}+\log^{3}K(\log(4K)+\log_{2}\frac{1}{\Delta_{j,0}^{2}})H(j)\right\}\right)&\nu\text{ is positive}\\ O\left((\log K)H_{1}^{\text{neg}}\log\frac{H_{1}^{\text{neg}}}{\delta}\right)&\nu\text{ is negative}\end{cases}$
Combining Theorems 2 and 3, we conclude that Algorithm 1 is Strong $\delta$-PAC. Moreover, the gap between the lower bound in Theorem 1 and the upper bound in Theorem 3 is at most polynomial in $\log K$ and $\{\log(1/\Delta_{0,a}^{2})\}_{a=1}^{m}$. This shows that PSEEB is nearly optimal for all 1-identification instances and identifies $\min_{1\leq j\leq m}\left\{\frac{\log(1/\delta)}{\Delta_{j,0}^{2}}+H(j)\right\}$ as the appropriate instance-dependent complexity measure on positive instances.
Compared to prior work, our results resolve an open problem posed in Li and Cheung (2025) by providing the first nearly optimal upper bound for instances with multiple qualified arms.
@@ 1.4 Literature Review on Other Related Works
If the threshold $\mu_{0}$ is viewed as an additional arm with deterministic reward, then 1-identification can be reduced to a standard Best-Arm Identification (BAI) problem Even-Dar et al. (2002); Gabillon et al. (2012); Jamieson et al. (2013); Kalyanakrishnan et al. (2012); Karnin et al. (2013); Jamieson and Nowak (2014); Chen and Li (2015). However, this reduction is known to be inefficient. Both asymptotic analyses Degenne and Koolen (2019); Garivier and Kaufmann (2016) and non-asymptotic results Li and Cheung (2025); Kano et al. (2017) show that such a formulation incurs unnecessary sampling. Intuitively, 1-identification does not require high-confidence pairwise comparisons among all arms, and therefore does not necessitate the intensive sampling in BAI Algorithms.
Beyond 1-identification, Degenne and Koolen (2019) research on several pure-exploration variants with multiple correct answers. These include identifying all arms in $\{a:\mu_{a}\geq\mu_{1}-\epsilon\}$ Mason et al. (2020), or identifying a single arm from this set Even-Dar et al. (2002); Gabillon et al. (2012); Kalyanakrishnan et al. (2012); Kaufmann and Kalyanakrishnan (2013); Katz-Samuels and Jamieson (2020). As observed in Katz-Samuels and Jamieson (2020), when the gap $\Delta_{0,1}$ is known and $\mu_{1}>\mu_{0}$, the answer sets of $\epsilon$-good arm identification and 1-identification coincide by setting $\epsilon=\Delta_{0,1}$, allowing the same sampling and recommendation rules. When the gaps are unknown, however, the precise relationship between these problems remains unclear. In addition, thresholding bandits Locatelli et al. (2016); Mukherjee et al. (2017) aim to identify arms with mean value above $\mu_{0}$ under a fixed budget. The correct output might not be unique under their proposed problem setting. Kano et al. (2017) also target at identifying all the arms with mean value above $\mu_{0}$, but they focus on the fixed confidence setting. Since Kano et al. (2017) provide an upper bound for each outputting time index, their result can be generalized to 1-identification.
Finally, our algorithmic design builds on the idea of running multiple algorithm copies in parallel, which has appeared in prior work. In particular, Chen and Li (2015); Chen et al. (2017); Katz-Samuels and Jamieson (2020) employ a master procedure to manage multiple executions of a base algorithm. The approach of Katz-Samuels and Jamieson (2020) allocates $2^{\ell}$ rounds to the execution indexed by $\ell$, while Chen and Li (2015); Chen et al. (2017) schedule parallel executions according to a global round index $t\in\mathbb{N}$, activating an execution when its index divides $t$. In contrast, our method adopts a simple round-robin scheduling scheme, which enables a balanced execution of parallel algorithm instances.
2 Lower Bound of Sample Complexity, for Positive Instances
@@
In this section, we continue to work with a reward vector $\{\mu_{a}\}_{a=1}^{K}$ satisfying $\mu_{1}\geq\mu_{2}\geq\cdots\geq\mu_{m}>\mu_{0}\geq\mu_{m+1}\geq\cdots\geq\mu_{K}$ and consider the instance $\nu$ defined by $r_{a}^{\nu}=\mu_{a}$. Throughout the lower-bound analysis, we assume unit-variance Gaussian rewards. For any instance $\nu^{\prime}$ and any permutation $\sigma:[K]\to[K]$ (a bijection), we define the permuted instance $\sigma(\nu^{\prime})$ by $r^{\sigma(\nu^{\prime})}_{\sigma(a)}\coloneqq r_{a}^{\nu^{\prime}},\forall a\in[K]$. We begin by introducing a symmetry notion that allows us to restrict attention to algorithms that do not prefer specific arm labels.
Definition 3 (Symmetric algorithm).
@@
An algorithm $\widetilde{\textnormal{alg}}$ is symmetric if for any instance $\nu$ with reward vector $\{\mu_{a}\}_{a=1}^{K}$, any sequence $\{a_{t}\}_{t=1}^{T}\in[K]^{T}$, any integer $T\geq 1$, and any two permutations $\sigma^{\prime},\sigma^{\prime\prime}:[K]\to[K]$, it holds that
$\Pr_{\sigma^{\prime}(\nu),\,\widetilde{\textnormal{alg}}}\!\bigl(\forall t\in[T],\,A_{t}=\sigma^{\prime}(a_{t}),\,\tau>T\bigr)=\Pr_{\sigma^{\prime\prime}(\nu),\,\widetilde{\textnormal{alg}}}\!\bigl(\forall t\in[T],\,A_{t}=\sigma^{\prime\prime}(a_{t}),\,\tau>T\bigr).$
@@
Intuitively, under a symmetric algorithm $\widetilde{\text{alg}}$, the reward distribution of the pulled arm $A_{t}$ at time $t$ is invariant under any permutation on the arms’ indexes. In particular, we have $\mathbb{E}_{\sigma^{\prime}(\nu),\,\widetilde{\textnormal{alg}}}[\tau]=\mathbb{E}_{\sigma^{\prime\prime}(\nu),\,\widetilde{\textnormal{alg}}}[\tau],\text{for permutations }\sigma^{\prime},\sigma^{\prime\prime}$. A similar definition also appears in (Simchowitz et al., 2017, Definition 12), which further requires the invariance of output distribution; this extra requirement is not needed here. The following lemma shows that it suffices to consider symmetric algorithms for lower bounds.
@@
Let $S$ denote the set of all permutations of $[K]$. For any $\delta$-PAC algorithm alg, there exists a $\delta$-PAC symmetric algorithm $\widetilde{\textnormal{alg}}$ such that for every instance $\nu^{\prime}$, $\max\limits_{\sigma\in S}\mathbb{E}_{\textnormal{alg},\,\sigma(\nu^{\prime})}[\tau]\geq\mathbb{E}_{\widetilde{\textnormal{alg}},\,\nu^{\prime}}[\tau]$.
The proof of Lemma 4 is in Appendix B.1. It is obtained by randomizing a permutation of arm labels prior to executing alg. Given Lemma 4, to prove Theorem 1, it suffices to establish the following lower bound for symmetric algorithms.
@@
For any symmetric and $\delta$-PAC alg and any reward vector $\{\mu_{a}\}_{a=1}^{K}\in[0,1]^{K}$ satisfying that $\mu_{1}\geq\cdots\geq\mu_{m}>\mu_{0}\geq\mu_{m+1}\geq\cdots\geq\mu_{K}$, any tolerance level $\delta$ such that $\delta<\min\{10^{-8},\frac{1}{64m^{2}}\}$, define a positive instance $\nu$ with $r_{a}^{\nu}=\mu_{a},a\in[K]$ with unit-variance Gaussian reward. We have $\mathbb{E}_{\text{alg},\nu}\tau\geq\Omega\left(\min\limits_{1\leq j\leq m}\left\{\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}+\frac{H(j)}{\log^{2}(m+1)}\right\}\right)$.
The assumption $\delta<\min\{10^{-8},1/(64m^{2})\}$ is to ensure $\delta\cdot m^{2}$ is upper bounded by an absolute constant. The upper bound eliminates the possibility of being $\delta$-PAC by outputting a uniformly chosen arm without any arm pulls.
@@
We outline the main steps; the full proof is in Appendix B.2. The main idea is to establish Lemma 6,
@@
Assume $\delta<\min\{10^{-8},1/(64m^{2})\}$, for any symmetric and $\delta$-PAC alg, $\mathbb{E}_{\nu,\text{alg}}\tau$ is lower bounded by the $\frac{1}{3200}$ times optimal value of the following optimization problem
@@
Given Lemma 6, we proceed to prove the Theorem 1 as follows. We first prove the above optimization problem (2) is convex. Then, we turn to construct an explicit feasible dual solution, and prove the objective value of this feasible solution is lower bounded by $\Omega\left(\min\limits_{1\leq j\leq m}\left\{\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}+\frac{H(j)}{\log^{2}(m+1)}\right\}\right)$. By the strong duality, we complete the proof of Theorem 5.
Denote event $\mathcal{E}_{j}=\left\{N_{j}(\tau)\geq\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}},\hat{a}=j\right\}$. To prove Lemma 6, we show taking $v=3200\mathbb{E}_{\nu}\tau$, $p_{j}=\Pr_{\nu}(\mathcal{E}_{j})$ for $j\in[m]$ is a feasible solution to (2). The argument repeatedly applies the transportation (change-of-measure) inequality (Kaufmann et al., 2016, Lemma 18) to relate $\Pr_{\nu}(\mathcal{E}_{j})$ to probabilities under suitable alternative instances. We prove the feasibility of $v=3200\mathbb{E}_{\nu}\tau$, $p_{j}=\Pr_{\nu}(\mathcal{E}_{j})$ by going through the constraint (C1), (C2), (C3), (C4) (C5).
For (C1), (C2), they are evident. As the expected pulling times must be non-negative and the probability must be in the interval $[0,1]$.
@@
$3200\mathbb{E}\tau\geq\mathbb{E}\tau\geq\sum_{j=1}^{m}\mathbb{E}\tau\mathds{1}(\mathcal{E}_{j})\geq\sum_{j=1}^{m}\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}}\Pr_{\nu}(\mathcal{E}_{j}).$
For (C4), we start by proving
$\Pr_{\nu}(\tau<+\infty,\hat{a}\in[m])\geq 1-5\delta,$ (3)
which must hold regardless of whether $\{a:\mu_{a}=\mu_{0}\}$ is empty. The main idea is to utilize the assumption that the algorithm is $\delta$-PAC.
Then, we are ready to prove (C4) by contradiction. Denote $\bar{\mathcal{E}_{j}}=\{N_{j}(\tau)<\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}},\hat{a}=j\}$. If $\sum_{j=1}^{m}\Pr_{\nu}(\mathcal{E}_{j})<\frac{1}{2}$, together with (3), we can find a $j_{0}\in[m]$ such that
@@
We consider an alternative instance $\nu^{\prime}_{j_{0}}(\epsilon)$ with $r_{a^{\prime}}^{\nu^{\prime}_{j_{0}}}=\begin{cases}\mu_{0}-\epsilon&a^{\prime}=j_{0}\\ r_{a^{\prime}}^{\nu}&else\end{cases}$. The $\delta$-PAC requirement of algorithm means that
@@
Combining (5),(4), we have $\log\frac{\Pr_{\nu}\left(\bar{\mathcal{E}}_{j_{0}}\right)}{\Pr_{\nu^{\prime}_{j_{0}}(\epsilon)}(\bar{\mathcal{E}}_{j_{0}})}\geq\frac{1}{2}\log\frac{1}{\delta}$ holds for all $\epsilon>0$. Meanwhile, Lemma 9, which utilizes the Transportation Equality (Kaufmann et al., 2016, Lemma 18), suggests that
@@
for each $j\in[m],a\in[K],j\neq a$, where $T_{j}(a)=\lceil\frac{1}{200\max\{\Delta_{j,0}^{2},\Delta_{a,j}^{2}\}\Big(1+\log\frac{1}{\Pr_{\nu}(\mathcal{E}_{j})}\Big)}\rceil$. Given Lemma 12 holds, sum up (7) for each $a\in[K]$ and use the fact that $\frac{1}{\max\{\Delta_{j,0}^{2},\Delta_{a,j}^{2}\}}\geq\frac{1}{4\max\{\Delta_{a,0}^{2},\Delta_{a,j}^{2}\}}$, we prove (C5). Then, the remaining work is to prove (7). In the following, we present the sketch idea, which highly relies on the definition of $T_{j}(a)$.
For $n\in\mathbb{N}$, we define $\tau_{n}^{j}(a)$ by looking for an time index $t$, such that $N_{j}(t)+N_{a}(t)=n,A_{t+1}\in\{j,a\}$. Such a time index $t$ is unique and $t<\tau$, if it exists. We take $\tau_{n}^{j}(a)=t$ if $t$ exists, or $\tau_{n}^{j}(a)=+\infty$ if $t$ doesn’t exist. Since $A_{t+1}$ is fully determined by the history up to time index $t$, for $t\in\mathbb{N}$. $\tau_{n}^{j}(a)$ is a stopping time regrading the filtration $\{\sigma(H(t))\}_{t\geq 1}$.
Denote $\tilde{\mathcal{E}}^{aj}(n)=\{\tau_{n}^{j}(a)<\tau\}$, $\tilde{\mathcal{E}}^{aj}_{j}(n)=\{A_{\tau_{n}^{j}(a)+1}=j,\tau_{n}^{j}(a)<\tau\}$, $\tilde{\mathcal{E}}^{aj}_{a}(n)=\{A_{\tau_{n}^{j}(a)+1}=a,\tau_{n}^{j}(a)<\tau\}$. We prove two inequalities.
$\Pr_{\nu}(\tilde{\mathcal{E}}^{aj}(n))\geq\Pr_{\nu}(\mathcal{E}_{j}),\forall n\leq T_{j}(a)-1.$ (8)
$\Pr_{\nu}(\tilde{\mathcal{E}}^{aj}_{a}(n))\geq\frac{1}{4}\Pr_{\nu}(\tilde{\mathcal{E}}^{aj}(n)),\forall n\leq T_{j}(a)-1.$ (9)
Given (8) and (9), from the fact that we only pull arm a at $\tau_{n}^{j}(a)+1$, we can conclude
[…]
Thus, to prove (7), we suffice to show (8) (9) are both correct.